# SonicMaster: Towards Controllable All-in-One Music Restoration and Mastering

### v2

@@ A.4 Spectrogram examples
Figure 6: Effect of clarity degradation and restoration on spectrograms. The treble frequencies are supressed in the degraded input sample, and then restored with SonicMaster. Prompt: ”Make the audio clearer and more intelligible.”
@@
Figure 7: Effect of clipping degradation and related restoration. Drum hits clip in the degraded audio, but are restored in the SonicMaster’s output without distortion. Prompt: ”Clean up the harshness in the signal.”
@@
Figure 8: Another example of the effect of clipping and its restoration. The degraded input shows signs of distortion with visible spectral content above 15 kHz. This distortion is suppressed by SonicMaster. Prompt: ”Clean up the noisiness in the audio.”
@@
A.5 Ablation on ODE Solvers
We evaluated Euler solvers with $1$, $10$ (baseline), and $100$ steps, plus a $10$-step 4th order Runge–Kutta (Dormand & Prince, 1980) solver. Tables A.4, A.4, and A.4 outline the results and highlight the trade-off across degradation categories. Euler-1 matches the baseline overall but is weaker on Boom, Microphone, Clip, all Reverb subtasks, and shows higher KL. Euler-100 boosts Reverb and Punch yet lowers every EQ score versus the 1-/10-step runs. Runge–Kutta-10 equals Euler-10 on most metrics and tops Clip, but its inference is significantly slower.
A.6 Prompts for each degradation type

### v3

@@ A.6 Spectrogram examples
Figure 9: Effect of clarity degradation and restoration on spectrograms. The treble frequencies are supressed in the degraded input sample, and then restored with SonicMaster. Prompt: “Make the audio clearer and more intelligible.”
@@
Figure 10: Effect of clipping degradation and related restoration. Drum hits clip in the degraded audio, showing as wideband spectral peaks, but are restored in the SonicMaster’s output without distortion. Prompt: “Clean up the harshness in the signal.”
@@
Figure 11: Another example of the effect of clipping and its restoration. The degraded input shows signs of distortion with visible increase in wideband spectral content at the parts of waveform clipping. This distortion is suppressed by SonicMaster. Prompt: “Clean up the noisiness in the audio.”
@@
A.7 Ablation on ODE solvers, model size, and conditioning
We evaluated Euler solvers with $1$, $10$ (baseline), and $100$ steps, plus a $10$-step 4th order Runge–Kutta (dormand1980family) solver. Tables 8, 9, and 10 outline the results and highlight the trade-off across degradation categories. Euler-1 matches the baseline overall but is weaker on Boom, Microphone, Clip, all Reverb subtasks, and shows higher KL. Euler-100 boosts Reverb and Punch yet lowers every EQ score versus the 1-/10-step runs. Runge–Kutta-10 equals Euler-10 on most metrics and tops Clip, but its inference is significantly slower.
We further performed a scaling analysis of the SonicMaster model. The results in Tables 8, 9, 10, show that SonicMasterSmall performs comparably with SonicMasterLarge in all metrics, but slightly worse in Reverb, Clip, and Stereo. The medium variant, SonicMasterMedium (4MM-DiT/12DiT), performs slightly better than the small model SonicMasterSmall overall. It also performs comparably to the large model SonicMasterLarge, outperforming it in Boom, or Compression, but still lacking behind in Clip. SonicMasterMedium (6MM-DiT/6DiT) performs the worst out of all variants across all metrics, suggesting a non-optimal ratio of MM-DiT to DiT blocks.
Regarding the audio condition and its duration, we evaluated SonicMasterLarge with three different conditioning lengths (5s, 10s, 15s). The performance across configurations was found to be comparable (Tables 8, 9, 10). For our default model version, we chose the 10-second setting as it balances computational efficiency with temporal overlap for long-form processing. The variant that uses audio condition through the pooling layers during inference scored comparably to the default setup, however, we can observe improvement in Clip and Volume (Table 9). The model trained without audio conditioning performs similarly across the board, scoring the best in Boom (0.0658, Table 8), but shows a clear drop in Clip performance (2.055 vs 1.506, see Table 9), which highlights the importance of this condition for this reconstruction task.
A.8 Prompts for each degradation type